OpenStudy (rmpjingwei):

√(X+23)+√(X-12)=7 Why X is 13? Show Steps please help!

OpenStudy (anonymous):

\[\sqrt{x+23}+\sqrt{x-12}=7~~...(1)\] \[\left( x+23 \right)-\left( x-12 \right)=x+23-x+12=35\] \[\frac{ \left( x+23 \right)-\left( x-12 \right) }{ \sqrt{x+23}+\sqrt{x-12} }=\frac{ \left( \sqrt{x+23} \right)^2-\left( \sqrt{x-12} \right)^2 }{\sqrt{x+23}+\sqrt{x-12} }=\frac{ 35 }{7 }=5\] \[\frac{ \left( \sqrt{x+23}+\sqrt{x-12} \right)\left( \sqrt{x+23}-\sqrt{x-12} \right) }{( \sqrt{x+23}+\sqrt{x-12}) }=5\] \[\sqrt{x+23}-\sqrt{x-12}=5 ~~...(2)\] adding (1) and (2) \[2\sqrt{x+23}=7+5=12,\sqrt{x+23}=6\] squaring both sides x+23=36 x=?

OpenStudy (rmpjingwei):

thanks alot

OpenStudy (anonymous):

yw